OpenStudy (anonymous):

An hour-long call to St. Louis from Chicago at the normal day rate-costs $12. If Amanda makes an hour-long call between 11 PM and 8 AM, when the telephone company gives a 40% discount, how much will she pay for the call? $4.80 $7 $7.20 $8.75 $16.80

OpenStudy (ghazi):

she will pay 12-(discount of that duration)=12-(0.4*12)=12-4.8=$ 7.2

OpenStudy (anonymous):

u're the boss, bro!

OpenStudy (ghazi):

:) ....thanks....i am sure you are learning with me :)

OpenStudy (anonymous):

of course. i ask the questions that give me the hardest time.

OpenStudy (ghazi):

cool..